How many years old is a person if her age now is equal to her age in 3 years time * 3, less 3 times her age 3 years ago?
Answer by ewatrrr(24785)


Hi
Let x represent her age NOW
Question states***Write as we Read
x = 3(x+3) - 3(x-3)
solving for x
x = 3x+9 - 3x + 9
x = 18, her age NOW
CHECKING our Answer***
18 = 3*21 - 3* 15 = 63 - 45